AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Sat Apr 5 00:42:59 2025
Author
*
:
Subject
*
:
<p>10:30 Found ADC#4 on aida 1 and 4 are suddenly noisy. See attach 1-2. Checkloaded multiple times - no effect.</p> <p>Leakage current overnight was stable, but a clear transient can be observed yesterday around the times of the DSSD motion test. We did not observe this in the cave looking at the CAEN module - it might have been too fast.</p> <p>Now reducing bias to 50 V and will check differences later. FEE temperatures OK.</p> <p> </p> <p>10:50 Started DAQ. Rates are much higher in n+n strips, but p+n strips appear unaffected. See attach 4-5.</p> <p>Stopped DAQ. Will leave 50V until after lunch.</p> <p> </p> <p>11:27 Checked motor control status after yesterday's setup (see previous ELOG entry). See attach 6.</p> <p>IN counters are respectively: read, increases, latch (left), latch (right)</p> <p>Read is the number of pulses sent with the instruction to move. Increases is the number of times the instruction was considered. Latch on/off indicates the 1 s time interval during which commands sent to a motor arm (left or right) are considered double-firing or mistake, and ignored.</p> <p>The number of increases is the same for move in/out - good!</p> <p>The number of increases for move out is 1 unit less than move in - one double pulse / double command must have been received overnight and ignored.</p> <p>Note torque on right arm is 2 (arbitrary units) which is unusually high and may indicate some tension in the kapton wires. No apparent effect on DSSD2,4 on that arm. Will not move for now.</p> <p> </p> <p>13:24 FEE temperatures OK, leakage currents decreasing slighly.</p> <p>13:41 Slowly increased bias to 100V. Interestingly it is now DSSD#4 (Ch#3) that displays the highest leakage current. Attach 7</p> <p>13:45 Stats & Rates look rather similar to earlier this morning at 150V, but aida01 and 04 are quieter. Attach 8-9. Will leave at 100V for a while longer before moving to 150V.</p> <p> </p> <p>15:20 Set bias to 150V</p> <p>15:50 Nikos testing target. Started short run. Stats & Rates attached (10-11). Same as yesterday, aida01 and 04 ADC#4 no longer noisy.</p> <p>V-I on attach 12-13. Legends unreliable!</p> <p> </p> <p>17:00 Nikos completed target test. YR09 pressure around 3E-10 mbar with around 5E11 atoms/cm2 He target. Target not quite stable (10-20%), some optimisation required, but pressure holds at 3E-10 mbar for several minutes.</p> <p>Beam on target will increase ring pressure. Can't test for now. Target tests without beam complete. Need to log all the data remotely. No logging available at present.</p> <p>Leaving bias at 150V overnight. No run.</p> <p> </p> <p> </p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown